# FBSDiff: Plug-and-Play Frequency Band Substitution of Diffusion Features for Highly Controllable Text-Driven Image Translation

### v2

@@ I Introduction
As a typical application of the booming multimodal technology, text-driven I2I translation is an appealing computer vision problem that aims to translate a reference image as per a text prompt. It extends text-to-image (T2I) synthesis to more controllability by controlling T2I generation result with a reference image. Since the advent of CLIP [1] bridging vision and language through large-scale contrastive pre-training, attempts have been made to instruct image manipulation with text by combining CLIP with generative models. VQGAN-CLIP [2] pioneers text-driven image translation by optimizing VQGAN [3] latent image embedding with CLIP text-image similarity loss. DiffusionCLIP [4] fine-tunes diffusion model [5] under the same CLIP loss to manipulate an image as per a text. DiffuseIT [6] combines VIT-based structure loss [7] and CLIP-based semantic loss to guide diffusion model’s reverse sampling process via manifold constrained gradient [8], synthesizing translated image that complies with the target text while maintaining the structure of the reference image. However, these methods are not competitive in generation visual quality due to the limited model capacity of backbone generative model as well as the inherent unstability caused by online fine-tuning or optimization process.
To promote image translation visual quality, efforts have been made to train large models on massive data. InstructPix2Pix [9] employs GPT-3 [10] and Stable Diffusion [11] to synthesize huge amounts of paired training data, based on which trains a supervised text-driven I2I mapping for general image manipulation task. Design Booster [12] trains a latent diffusion model [11] conditioned on both text embedding and image embedding, realizing layout-preserved text-driven I2I translation. Nevertheless, these methods are computationally intensive in training large models from scratch and less efficient in collecting immense training data.
@@
The former type of fine-tuning-based methods represented by SINE [13] and Imagic [14] fine-tune the pre-trained T2I diffusion model to reconstruct an input reference image before manipulating it with a target text. These methods require separate fine-tuning of the entire diffusion model for each time of image manipulation, which is less efficient and prone to underfitting or overfitting to the reference image.
The latter type of inversion-based methods invert reference image into diffusion model’s Gaussian noise space and then generate the translated image via the reverse sampling process guided by the target text. A pivotal challenge of this pipeline is that the sampling trajectory may severely deviate from the inversion trajectory due to the error accumulation caused by the classifier-free guidance technique [15], which severely impairs the correlation between the reference image and the translated image. To remedy this issue, Null-text Inversion [16] optimizes the unconditional null-text embedding to calibrate the sampling trajectory step by step. Prompt Tuning Inversion [17] proposes to minimize trajectory divergence with an optimization to encode the reference image into a learnable prompt embedding. Similarly, StyleDiffusion [18] opts to optimize the “value” embedding of the cross-attention layer as the visual encoding of the reference image. Pix2Pix-zero [19] penalizes trajectory deviation by matching cross-attention maps between the two trajectories with least-square loss. These methods apply per-step online optimization to calibrate the whole sampling trajectory, introducing additional computational cost and time overhead. Moreover, most of these methods adopt the cross-attention control technique introduced in Prompt-to-Prompt [20] for image structure preservation. This makes them rely on a paired source text of the reference image, which is not flexible or even available in most cases. Plug-and-Play (PAP) [21] proposes to leverage feature maps and self-attention maps extracted from internal layers of the denoising U-Net to maintain image structure, realizing optimization-free text-driven I2I translation. However, the algorithm is sensitive to specific layer selection, and the feature extraction process is also time-consuming.
@@ II-A Diffusion Model
Since the advent of DDPM [5], diffusion model has soon dominated the family of generative models [22]. Afterwards, much progress have been made to improve diffusion model in both methodology and application. DDIM [23] and its variants [24, 24] accelerate diffusion model sampling process to tens of times with only marginal drop in generation quality, promoting its practicability dramatically. Palette [25] extends diffusion model from unconditional image generation to the realm of conditional image synthesis, opening the door of diffusion-based image-to-image translation. With the advancement of multimodal technology, large-scale T2I diffusion models [26, 27, 28] are proposed to generate high-resolution images with open-domain text prompts, bringing content creation to an unprecedented level. To lower computational overhead of large-scale T2I model, Latent Diffusion Model (LDM) [11] proposes to transfer diffusion model from high-dimension pixel space to low-dimensional feature space, contributing the most widely used architecture in AIGC industry. To introduce more controllability to T2I synthesis, ControlNet [29] and T2i-adapter [30] add spatial control to T2I diffusion models by training a control module of the denoising U-Net conditioned on certain image priors (e.g., canny edges, depth maps, human key points, etc.). SDXL [31] and DiTs [32] propose Transformer [33] based backbone denoising network, improving T2I diffusion model to larger capacity. Up to now, diffusion model has been applied to a wide variety of vision fields such as image super-resolution [34], image inpainting [35], image colorization [36], semantic segmentation [37], point cloud generation [38], video synthesis [39], 3D reconstruction [40], etc, and is still making rapid progress in theory and potential applications.
@@ II-B Computer Vision in Frequency Perspective
Deep neural network models are mostly applied to tackle vision tasks in spatial or temporal domain, some research reveals that model performance can also be boosted from frequency domain. For example, Ghosh et al. [41] introduce DCT to convolutional neural network for image classification, accelerating network convergence speed. Xie et al. [42] propose a frequency-aware dynamic network for lightweight image super-resolution. Cai et al. [43] impose Fourier frequency spectrum consistency to image translation tasks, achieving better identity preservation ability. FreeU [44] improves T2I generation quality by selectively enhancing or depressing different frequency components of diffusion features inside the denoising U-Net model. ILVR [45] proposes to fuse low-frequency information of the forward diffusion process into the reverse sampling process for conditioned image synthesis. Our method differs with ILVR in that ILVR simulates low-pass filtering with simple feature downsampling and upsampling and performs information fusion in the spatial domain, while our method explicitly extracts and transplants frequency bands of diffusion features in the DCT domain. FCDiffusion [46] shares similar spirit of frequency-based control of T2I diffusion model with our method. However, FCDiffusion relies on training multiple frequency control branches to realize versatile control effects, while our method achieves versatility and high controllability in both guiding factor and guiding intensity of the reference image in a training-free and plug-and-play manner.
@@ III-A Overall Architecture
where $\bar{\alpha}_{t}$} are schedule parameters that follows the same setting as DDPM [5], $\epsilon_{\theta}$ is the denoising U-Net of the pre-trained LDM. The Gaussian noise $z_{T_{inv}}$ obtained after the $T_{inv}$-step DDIM inversion is directly used as the initial noise feature of the subsequent reconstruction trajectory, which is a $T$-step DDIM sampling process that reconstructs $\hat{z}_{0}\approx z_{0}$ from the inverted noise feature $\hat{z}_{T}=z_{T_{inv}}$:
@@
Meanwhile, an equal-length sampling trajectory is applied in parallel with the reconstruction trajectory for T2I synthesis. The sampling trajectory is also a $T$-step DDIM sampling process that progressively denoises a randomly initialized Gaussian noise feature $\tilde{x}_{T}\sim\mathcal{N}(0,I)$ into $\tilde{x}_{0}$ conditioned on the text embedding $v$ of the target text prompt. To amplify the effect of text guidance, we employ classifier-free guidance technique [15] by interpolating conditional (target text) and unconditional (null text) noise prediction at each time step with a guidance scale $\omega$ along the sampling process:
@@ III-B Frequency Band Substitution Layer
[…]
@@ III-C Implementation Details
11:  for $t=\lambda T$ to $1$ do
12:     compute $\tilde{z}_{t-1}$ from $\tilde{z}$ via Eq. 4;
@@
We use the pre-trained Stable Diffusion v1.5 as the backbone diffusion model and set the classifier-free guidance scale $\omega=7.5$. We use 1000-step DDIM inversion to ensure high-quality reconstruction, i.e., $T_{inv}$=1000, and use 50-step DDIM sampling for both the reconstruction and sampling trajectory, i.e., $T$=50. Along the sampling trajectory, we allocate 55% time steps to the calibration phase and the remaining 45% steps for the non-calibration phase, i.e., $\lambda$=0.45. For the default DCT masking thresholds used in the FBS layer, we set $th_{lp}$=80 for low-frequency band substitution (low-FBS); $th_{hp}$=5 for high-frequency band substitution (high-FBS); $th_{mp1}$=5, $th_{mp2}$=80 for mid-frequency band substitution (mid-FBS). The complete algorithm of FBSDiff is presented in Alg. 1.
@@ IV-A Qualitative Results
We qualitatively compare our method with SOTA text-driven I2I translation methods including Plug-and-Play (PAP) [21], Null-text Inversion (Null-text) [16], Pix2Pix-zero [19], InstructPix2Pix (InsPix2Pix) [9], Prompt Tuning Inversion (PT-inversion) [17], StyleDiffusion [18], and VQGAN-CLIP (VQCLIP) [2], results are displayed in Fig. 7. The top panel of Fig. 7 shows that our method with low-FBS achieves better appearance consistency between the reference image and the translated result than related approaches, and is thus better suited to image creation scenario which favors inheriting the appearance and style from an existing image. The bottom panel of Fig. 7 shows that existing SOTA text-driven I2I methods struggle at producing I2I results with large appearance change from the reference images, while our method with high-FBS excels in generating images with significantly different appearance, and is thus more suitable to image creation scenario where appearance divergence is pursued. Among the compared approaches, our method is the only one that enables flexible control over different guiding factors of the reference image, and is also the only approach that simultaneously dispenses with model training, fine-tuning, online optimization, and attention modulations.
An advantage of our approach over related methods is sampling diversity. As displayed in Fig. 6, our FBSDiff can produce diverse text-guided I2I results by randomly sampling $\tilde{x}_{T}$ from isotropic Gaussian distribution, while other inversion-based methods [16, 21, 17, 19, 18] lack such sampling diversity due to directly initializing $\tilde{x}_{T}$ with the inverted feature embedding of the reference image.
@@
PAP [21] 0.954 0.272 20.440 0.287 6.590 0.956 28.337 0.279 6.458
Null-text [16] 0.948 0.247 17.546 0.276 6.505 0.952 22.545 0.270 6.402
Pix2Pix-zero [19] 0.951 0.243 16.875 0.262 6.484 0.953 21.240 0.258 6.344
InsPix2Pix [9] 0.958 0.266 23.373 0.258 6.269 0.965 30.804 0.264 6.196
PT-inversion [17] 0.947 0.248 21.667 0.271 6.481 0.948 24.367 0.267 6.285
StyleDiffusion [18] 0.944 0.251 22.484 0.267 6.477 0.947 25.166 0.260 6.267
@@ IV-B Ablation Study
To verify the rationality and effectiveness of our proposed method, we also explore and compare with other designs of frequency band substitution, including substituting the frequency band only once at $\lambda T$ time step rather than along the whole calibration phase (which we denote as Once Substitution), and substituting the full DCT spectrum rather other only a partial frequency band of it (which we refer to as Full Substitution).
@@ IV-C Quantitative Evaluations
For quantitative method evaluation, we separately evaluate methods on the text-driven I2I translation task pursuing image appearance consistency and the task pursuing image appearance divergence. For the former task, we assess models’ appearance and layout preservation ability by measuring Structure Similarity ($\uparrow$), Perceptual Similarity ($\uparrow$), and Style Distance ($\downarrow$) between the reference image and the translated image pair. For the latter task, we assess models’ contour preservation and appearance alteration capabilities by measuring Structure Similarity ($\uparrow$) and Style Distance ($\uparrow$) between I2I translation pairs. For Structure Similarity measurement, we use DINO-ViT self-similarity distance [7] as the metric for Structure Distance between two images, and define Structure Similarity as 1 - Structure Distance. We use LPIPS [47] metric to measure Perceptual Similarity, and use AdaIN style loss [48] to measure Style Distance between I2I pairs. Besides, CLIP Similarity ($\uparrow$) metric is used to measure semantic consistency between the target text prompt and the translated image, i.e., text fidelity of the I2I translation results. Finally, we evaluate Aesthetic Score ($\uparrow$) of the translated images via the pre-trained LAION Aesthetics Predictor V2 model.
@@
Null-text [16]
PAP [21]
Pix2Pix-zero [19]
InsPix2Pix [9]
PT-inversion [17]
StyleDiffusion [18]
VQCLIP [2]
DiffuseIT [6]
DiffusionCLIP [4]
Design Booster [12]
SINE [13]
Imagic [14]

### v6

@@ I Introduction
As a typical application of the booming multimodal technology, text-driven I2I translation is an appealing computer vision problem that aims to translate a reference image as per a text prompt. It extends text-to-image (T2I) synthesis to more controllability by controlling T2I generation result with a reference image. Since the advent of CLIP [radford2021learning] bridging vision and language through large-scale contrastive pre-training, attempts have been made to instruct image manipulation with text by combining CLIP with generative models. VQGAN-CLIP [crowson2022vqgan] pioneers text-driven image translation by optimizing VQGAN [esser2021taming] latent image embedding with CLIP text-image similarity loss. DiffusionCLIP [kim2022diffusionclip] fine-tunes diffusion model [ho2020denoising] under the same CLIP loss to manipulate an image as per a text. DiffuseIT [kwon2022diffusion] combines VIT-based structure loss [tumanyan2022splicing] and CLIP-based semantic loss to guide diffusion model’s reverse sampling process via manifold constrained gradient [chung2022improving], synthesizing translated image that complies with the target text while maintaining the structure of the reference image. However, these methods are not competitive in generation visual quality due to the limited model capacity of backbone generative model as well as the inherent unstability caused by online fine-tuning or optimization process.
To promote image translation visual quality, efforts have been made to train large models on massive data. InstructPix2Pix [brooks2023instructpix2pix] employs GPT-3 [brown2020language] and Stable Diffusion [rombach2022high] to synthesize huge amounts of paired training data, based on which trains a supervised text-driven I2I mapping for general image manipulation task. Design Booster [sun2023design] trains a latent diffusion model [rombach2022high] conditioned on both text embedding and image embedding, realizing layout-preserved text-driven I2I translation. Nevertheless, these methods are computationally intensive in training large models from scratch and less efficient in collecting immense training data.
@@
The former type of fine-tuning-based methods represented by SINE [zhang2023sine] and Imagic [kawar2023imagic] fine-tune the pre-trained T2I diffusion model to reconstruct an input reference image before manipulating it with a target text. These methods require separate fine-tuning of the entire diffusion model for each time of image manipulation, which is less efficient and prone to underfitting or overfitting to the reference image.
The latter type of inversion-based methods invert reference image into diffusion model’s Gaussian noise space and then generate the translated image via the reverse sampling process guided by the target text. A pivotal challenge of this pipeline is that the sampling trajectory may severely deviate from the inversion trajectory due to the error accumulation caused by the classifier-free guidance technique [ho2022classifier], which severely impairs the correlation between the reference image and the translated image. To remedy this issue, Null-text Inversion [mokady2023null] optimizes the unconditional null-text embedding to calibrate the sampling trajectory step by step. Prompt Tuning Inversion [dong2023prompt] proposes to minimize trajectory divergence with an optimization to encode the reference image into a learnable prompt embedding. Similarly, StyleDiffusion [li2023stylediffusion] opts to optimize the “value” embedding of the cross-attention layer as the visual encoding of the reference image. Pix2Pix-zero [parmar2023zero] penalizes trajectory deviation by matching cross-attention maps between the two trajectories with least-square loss. These methods apply per-step online optimization to calibrate the whole sampling trajectory, introducing additional computational cost and time overhead. Moreover, most of these methods adopt the cross-attention control technique introduced in Prompt-to-Prompt [hertz2022prompt] for image structure preservation. This makes them rely on a paired source text of the reference image, which is not flexible or even available in most cases. Plug-and-Play (PAP) [tumanyan2023plug] proposes to leverage feature maps and self-attention maps extracted from internal layers of the denoising U-Net to maintain image structure, realizing optimization-free text-driven I2I translation. However, the algorithm is sensitive to specific layer selection, and the feature extraction process is also time-consuming.
@@ II-A Diffusion Model
Since the advent of DDPM [ho2020denoising], diffusion model has soon dominated the family of generative models [dhariwal2021diffusion]. Afterwards, much progress have been made to improve diffusion model in both methodology and application. DDIM [song2020denoising] and its variants [lu2022dpm, lu2022dpm] accelerate diffusion model sampling process to tens of times with only marginal drop in generation quality, promoting its practicability dramatically. Palette [saharia2022palette] extends diffusion model from unconditional image generation to the realm of conditional image synthesis, opening the door of diffusion-based image-to-image translation. With the advancement of multimodal technology, large-scale T2I diffusion models [nichol2022glide, ramesh2022hierarchical, saharia2022photorealistic] are proposed to generate high-resolution images with open-domain text prompts, bringing content creation to an unprecedented level. To lower computational overhead of large-scale T2I model, Latent Diffusion Model (LDM) [rombach2022high] proposes to transfer diffusion model from high-dimension pixel space to low-dimensional feature space, contributing the most widely used architecture in AIGC industry. To introduce more controllability to T2I synthesis, ControlNet [zhang2023adding] and T2i-adapter [mou2024t2i] add spatial control to T2I diffusion models by training a control module of the denoising U-Net conditioned on certain image priors (e.g., canny edges, depth maps, human key points, etc.). SDXL [podell2023sdxl] and DiTs [peebles2023scalable] propose Transformer [vaswani2017attention] based backbone denoising network, improving T2I diffusion model to larger capacity. Up to now, diffusion model has been applied to a wide variety of vision fields such as image super-resolution [saharia2022image], image inpainting [lugmayrinpainting], image colorization [liang2024control], semantic segmentation [tan2023diffss], point cloud generation [luo2021diffusion], video synthesis [yu2023video], 3D reconstruction [anciukevivcius2023renderdiffusion], etc, and is still making rapid progress in theory and potential applications.
@@ II-B Computer Vision in Frequency Perspective
Deep neural network models are mostly applied to tackle vision tasks in spatial or temporal domain, some research reveals that model performance can also be boosted from frequency domain. For example, Ghosh et al. [ghosh2016deep] introduce DCT to convolutional neural network for image classification, accelerating network convergence speed. Xie et al. [xie2021learning] propose a frequency-aware dynamic network for lightweight image super-resolution. Cai et al. [cai2021frequency] impose Fourier frequency spectrum consistency to image translation tasks, achieving better identity preservation ability. FreeU [si2023freeu] improves T2I generation quality by selectively enhancing or depressing different frequency components of diffusion features inside the denoising U-Net model. ILVR [choi2021ilvr] proposes to fuse low-frequency information of the forward diffusion process into the reverse sampling process for conditioned image synthesis. Our method differs with ILVR in that ILVR simulates low-pass filtering with simple feature downsampling and upsampling and performs information fusion in the spatial domain, while our method explicitly extracts and transplants frequency bands of diffusion features in the DCT domain. FCDiffusion [gao2024frequency] shares similar spirit of frequency-based control of T2I diffusion model with our method. However, FCDiffusion relies on training multiple frequency control branches to realize versatile control effects, while our method achieves versatility and high controllability in both guiding factor and guiding intensity of the reference image in a training-free and plug-and-play manner.
@@ III-A Overall Architecture
where $\bar{\alpha}_{t}$} are schedule parameters that follows the same setting as DDPM [ho2020denoising], $\epsilon_{\theta}$ is the denoising U-Net of the pre-trained LDM. The Gaussian noise $z_{T_{inv}}$ obtained after the $T_{inv}$-step DDIM inversion is directly used as the initial noise feature of the subsequent reconstruction trajectory, which is a $T$-step DDIM sampling process that reconstructs $\hat{z}_{0}\approx z_{0}$ from the inverted noise feature $\hat{z}_{T}=z_{T_{inv}}$:
@@
Meanwhile, an equal-length sampling trajectory is applied in parallel with the reconstruction trajectory for T2I synthesis. The sampling trajectory is also a $T$-step DDIM sampling process that progressively denoises a randomly initialized Gaussian noise feature ${\tilde{z}}_{T}\sim\mathcal{N}(0,I)$ into ${\tilde{z}}_{0}$ conditioned on the text embedding $v$ of the target text prompt. To amplify the effect of text guidance, we employ classifier-free guidance technique [ho2022classifier] by interpolating conditional (target text) and unconditional (null text) noise prediction at each time step with a guidance scale $\omega$ along the sampling process:
@@ III-B Frequency Band Substitution Layer
$\left\{\begin{array}[]{lr}Mask_{lp}(x,y)=1\ \ if\ \ x+y\leq th_{lp}\ \ else\ \ 0,&\\ Mask_{hp}(x,y)=1\ \ if\ \ x+y>th_{hp}\ \ else\ \ 0,&\\ Mask_{mp}(x,y)=1\ \ if\ \ th_{mp1}<x+y\leq th_{mp2}\ \ else\ \ 0,\end{array}\right$.
@@ III-C Implementation Details
11:for $t=\lambda T$ to 11 do
12:  compute $\tilde{z}_{t-1}$ from $\tilde{z}_{t}$ via Eq. 4;
@@
We use the pre-trained Stable Diffusion v1.5 as the backbone diffusion model and set the classifier-free guidance scale $\omega=7.5$. We use 1000-step DDIM inversion to ensure high-quality reconstruction, i.e., $T_{inv}$=1000, and use 50-step DDIM sampling for both the reconstruction and sampling trajectory, i.e., TT=50. Along the sampling trajectory, we allocate 55% time steps to the calibration phase and the remaining 45% steps for the non-calibration phase, i.e., $\lambda$=0.45. For the default DCT masking thresholds used in the FBS layer, we set $th_{lp}$=80 for low-frequency band substitution (low-FBS); $th_{hp}$=5 for high-frequency band substitution (high-FBS); $th_{mp1}$=5, $th_{mp2}$=80 for mid-frequency band substitution (mid-FBS). The complete algorithm of FBSDiff is presented in Alg. 1.
@@ IV-A Qualitative Results
We qualitatively compare our method with SOTA text-driven I2I translation methods including Plug-and-Play (PAP) [tumanyan2023plug], Null-text Inversion (Null-text) [mokady2023null], Pix2Pix-zero [parmar2023zero], InstructPix2Pix (InsPix2Pix) [brooks2023instructpix2pix], Prompt Tuning Inversion (PT-inversion) [dong2023prompt], StyleDiffusion [li2023stylediffusion], and VQGAN-CLIP (VQCLIP) [crowson2022vqgan], results are displayed in Fig. 7. The top panel of Fig. 7 shows that our method with low-FBS achieves better appearance consistency between the reference image and the translated result than related approaches, and is thus better suited to image creation scenario which favors inheriting the appearance and style from an existing image. The bottom panel of Fig. 7 shows that existing SOTA text-driven I2I methods struggle at producing I2I results with large appearance change from the reference images, while our method with high-FBS excels in generating images with significantly different appearance, and is thus more suitable to image creation scenario where appearance divergence is pursued. Among the compared approaches, our method is the only one that enables flexible control over different guiding factors of the reference image, and is also the only approach that simultaneously dispenses with model training, fine-tuning, online optimization, and attention modulations.
An advantage of our approach over related methods is sampling diversity. As displayed in Fig. 6, our FBSDiff can produce diverse text-guided I2I results by randomly sampling $\tilde{z}_{T}$ from isotropic Gaussian distribution, while other inversion-based methods [mokady2023null, tumanyan2023plug, dong2023prompt, parmar2023zero, li2023stylediffusion] lack such sampling diversity due to directly initializing $\tilde{z}_{T}$ with the inverted feature embedding of the reference image.
@@
PAP [tumanyan2023plug] 0.954 0.272 20.440 0.287 6.590 0.956 28.337 0.279 6.458
Null-text [mokady2023null] 0.948 0.247 17.546 0.276 6.505 0.952 22.545 0.270 6.402
Pix2Pix-zero [parmar2023zero] 0.951 0.243 16.875 0.262 6.484 0.953 21.240 0.258 6.344
InsPix2Pix [brooks2023instructpix2pix] 0.958 0.266 23.373 0.258 6.269 0.965 30.804 0.264 6.196
PT-inversion [dong2023prompt] 0.947 0.248 21.667 0.271 6.481 0.948 24.367 0.267 6.285
StyleDiffusion [li2023stylediffusion] 0.944 0.251 22.484 0.267 6.477 0.947 25.166 0.260 6.267
@@ IV-B Ablation Study
To verify the rationality and effectiveness of our proposed method, we also explore and compare with other designs of frequency band substitution, including substituting the frequency band only once at $\lambda T$ time step rather than along the whole calibration phase (which we denote Once Substitution), and substituting the full DCT spectrum rather than a partial sub-band of it (which we refer to as Full Substitution).
@@ IV-C Quantitative Evaluations
For quantitative method evaluation, we separately evaluate methods on the text-driven I2I translation task pursuing image appearance consistency and the task pursuing image appearance divergence. For the former task, we assess models’ appearance and layout preservation ability by measuring Structure Similarity ($\uparrow$), Perceptual Similarity ($\uparrow$), and Style Distance ($\downarrow$) between the reference image and the translated image pair. For the latter task, we assess models’ contour preservation and appearance alteration capabilities by measuring Structure Similarity ($\uparrow$) and Style Distance ($\uparrow$) between I2I translation pairs. For Structure Similarity measurement, we use DINO-ViT self-similarity distance [tumanyan2022splicing] as the metric for Structure Distance between two images, and define Structure Similarity as 1 - Structure Distance. We use LPIPS [zhang2018unreasonable] metric to measure Perceptual Similarity, and use AdaIN style loss [huang2017arbitrary] to measure Style Distance between I2I pairs. Besides, CLIP Similarity ($\uparrow$) metric is used to measure semantic consistency between the target text prompt and the translated image, i.e., text fidelity of the I2I translation results. Finally, we evaluate Aesthetic Score ($\uparrow$) of the translated images via the pre-trained LAION Aesthetics Predictor V2 model.
@@
Null-text [mokady2023null]
PAP [tumanyan2023plug]
Pix2Pix-zero [parmar2023zero]
InsPix2Pix [brooks2023instructpix2pix]
PT-inversion [dong2023prompt]
StyleDiffusion [li2023stylediffusion]
VQCLIP [crowson2022vqgan]
DiffuseIT [kwon2022diffusion]
DiffusionCLIP [kim2022diffusionclip]
Design Booster [sun2023design]
SINE [zhang2023sine]
Imagic [kawar2023imagic]